A Bibliography of Matrix Analysis, Second Edition

Roger A. Horn University of Utah,

Department of Mathematics,

155 S 1400 E RM 233, Salt Lake City, UT 84112-0090, USA,

Tel: +1 801 585 6819, FAX: +1 801 581 4148,

e-mail: rhorn@math.utah.edu (Internet)

and

Nelson H. F. Beebe University of Utah,

Department of Mathematics, 110 LCB,

155 S 1400 E RM 233, Salt Lake City, UT 84112-0090, USA,

Tel: +1 801 581 5254, FAX: +1 801 581 4148,

e-mail: beebe@math.utah.edu,
 beebe@acm.org,
beebe@computer.org (Internet),
URL: https://www.math.utah.edu/~beebe/

13 May 2024 Version 1.06

Abstract

This is a bibliography of publications in the second edition of the book Matrix Analysis (2012).

Title word cross-reference

1 [ZF84].

2 [ZF86].

6 [Bar15].

978 [Bar15]. 978-0-521-54823-6 [Bar15].

= [HS96].

Advanced [OCV11, Rom92, Rom05, Rom08]. Age [Cri06]. Algebra [Axl96, Bap93, C^{+}02, Chi79, Chi95, Chi09, FF63, Fad59, Gol04, Gre67, Gre78, HS74, HK61, HK71, Hog07, Hou72, Kap03, Kow63, Kow67, Kow69, Kow70, Kow79, KM95, KM98, KM03, Lan87, Lax07, Mar73, Mar75, Mir55, Mir61, Mir63, Mir82, Mir90, ND69, ND77, ND88, NDG89, OCV11, Rom92, Rom05, Rom08, Zha09, HS96, Kow63, Kow67, Kow69, Kow70, Kow79, KM95, KM98, KM03]. Algebraic [Wil65, Wil88]. Algebras [BD71]. Analysis [Bel70, Bha97, Con90, GL06, HJ85, HJ91, HJ12, Hou64, Hou75, Hou06, Kre78, Rud53, Rud64, Rud76, Tan07a, Tan07b, Tan07c, Tod62, Var00, HS96, And87, Bar15]. Analytic [Bau85, Don74, WM67]. angewandte [ZF84, ZF86]. Anwendungen [ZF84, ZF86, Zur61, Zur64]. aplicada [NDG89]. Applications [BR97, Bar90, BL88, Fan59, Fie86, Gan59a, Gan05, GLR06, Gra81, Gra69, Gra83, Kre78, LT85, Lax07, Lay82, Lay92, Lay07, MO79, MOA11, PR71a, PR71b, Sen73b, Sen73a, Ser02, Zha05, Zur61, Zur64]. Applied [ND69, ND77, ND88, ZF84, ZF86]. Arthur [Cri06]. Assets [C^{+}02].

Basic [Zha11]. Beginning [Gol04]. Book [Hal67]. Booko [Hal82]. Bounds [Bha87].

Calculus [Gra81]. Cambridge [Bar15]. Canonical [TA32, TA52, TA61, TA05]. Cayley [Cri06]. Chains [Sen81, Sen06]. Challenging [Zha09]. Charles [And87, Bar15]. Circles [Var04]. Classes [Fie75]. Combinatorial [BR91, Kőn36]. Commutative [ST68]. Complement [Zha05]. Complex [Tan07a]. complexes [Kőn36]. Computation [Hig08]. Computational [FF63, Fad59]. Computations [GV96, Ste73]. Concrete [Chi79, Chi95, Chi09]. Continuation [Don74]. Contributions [Mui30]. Control [Bar75, Bar83, BC85]. Convex [Fan59, Lay82, Lay92, Lay07, Val64, Val76]. Course [Con90, Kap03, Ort87].

Darstellung [Zur50, Zur58]. Definite [Bha07]. Derivatives [Rog80]. Determinanten [Ait69]. Determinants [Ait39, Ait46, Ait48, Ait54, Ait56, Ait69, Ait83, Bow62a, Bow62b, Bro58, Mui23, Mui30, Mui60a, Mui60b, Tan07a, Tur50]. Development [Mui23, Mui60a]. Differential [Cou89, HS74, Tan07b, Tan07c]. Dimensional [GL06, Hal42, Hal58, Mar73, Mar75]. Done [Axl96]. Dynamic [BNS89]. Dynamical [HS74].

edition [Bar15]. Eigenvalue [Par80, Par98, Wil65, Wil88]. Eigenvalues [Bha87, Cha93]. Elements [BD71]. endlichen [Kőn36, Kőn50, Kőn54]. Engineers [Bar79, Tan07a, Tan07b, Tan07c, ZF84, ZF86]. Equations [Cou89, HS74, Tan07b, Tan07c]. etc [Bar15].

Facts [Ber09]. Finite [GL06, Hal42, Hal58, Kőn90, Mar73, Mar75, Kőn36, Kőn50, Kőn54]. Finite-Dimensional [GL06, Hal42, Hal58]. Form [GL06, OCV11]. Formulas [Ber09]. Foundations [ZF84]. Fourier [Tan07c]. Functional [Con90, Kre78]. Functions [Boa60, Boa72, Boa81, Boa96, Don74, Hig08].

Gems [C^{+}02]. Generalized [CM79, CM91, CM09, PR71a, PR71b]. Geometry [Kap03]. German [Kőn36, Kőn50, Kőn54, Ait69, CH24, CH37, CH42, CH68a, CH68b, GK60, Kow63, Kow67, Kow69, Kow70, Kow79, KM95, KM98, KM03, Zur61, Zur64, ZF84, ZF86]. Gersgorin [Var04]. Graduate [Gol04]. Graphen [Kőn36, Kőn50, Kőn54]. Graphs [Kőn90, Kőn36, Kőn50, Kőn54]. Grundlagen [ZF84].

Handbook [Hog07]. Helmut [HS96]. Higher [Chi79, Chi95, Chi09]. Hilbert [Hal67, Hal82]. Historical [Mui23, Mui60a]. History [Mui30]. Horn [And87, Bar15].

ihre [Zur61, Zur64, ZF84, ZF86]. II [BD73]. Indefinite [GLR83]. Inequalities [BB65, HLP34, HLP59, MM64, MM92, MO79, MOA11, Zha02]. Infinite [Kőn90, Kőn36, Kőn50, Kőn54]. Ingenieure [ZF84, ZF86, Zur50, Zur58]. Integral [New72]. Introduction [Bar75, BC85, Bel70, Bow62a, Bow62b, Bro58, Chi79, Chi95, Chi09, Fin60, Fin66, Fin78, Fin11, GK69, Gra69, Mir55, Mir61, Mir63, Mir82, Mir90, Sen73b, Sen73a, Sta69, Ste73, TA32, TA52, TA61, TA05]. Introductory [Kre78]. Invariant [GLR06]. Invariants [Tur50]. Inverse [PR71a, PR71b]. Inverses [CM79, CM91, CM09]. ISBN [Bar15]. Iterative [Var00].

James [Par06]. Jewish [Par06]. Johnson [Bar15, And87]. Joseph [Par06].

kernels [GK60]. kleine [GK60]. Know [Gol04]. kombinatorische [Kőn36]. Kronecker [Gra81].

Laplace [Tan07b]. Laureate [Cri06]. Least [LH74, LH95]. Lecture [WM67, Fan59]. Lectures [Hou72, Wed34, Wed64, Wed05]. lineal [NDG89]. Linear [Axl96, Bap93, Bar83, BS95, CM79, CM91, CM09, C^{+}02, Cul66, Cul72, Cul90, FF63, Fad59, Fin60, Fin66, Fin78, Fin11, GL06, GK69, Gol04, HS74, HK61, HK71, Hog07, Kap03, Kat66, Kat76, Kat84, Kat95, Kow63, Kow67, Kow69, Kow70, Kow79, KM95, KM98, KM03, Lan87, Lax07, Mir55, Mir61, Mir63, Mir82, Mir90, ND69, ND77, ND88, OCV11, Rom92, Rom05, Rom08, Sta69, Zha09, HS96]. Lineare [Kow63, Kow67, Kow69, Kow70, Kow79, KM95, KM98, KM03].

Majorization [MO79, MOA11]. Markov [Sen81, Sen06]. Mathematical [Ano60, Bar75, BC85, BP79, BP94, CH24, CH37, CH42, CH53, CH68a, CH68b, CH89, Cou89, HS96, Rud53, Rud64, Rud76, Tan07a, Tan07b, Tan07c]. Mathematician [Cri06, Par06]. Mathematicians [ZF84, ZF86]. Mathematics [Ber09, C^{+}02, Fie86, YG72, YG73, YG88a, YG88b]. Mathematiker [ZF84, ZF86]. Mathematische [HS96]. mathematischen [CH24, CH37, CH42, CH68a, CH68b]. Matrices [Ait39, Ait46, Ait48, Ait54, Ait56, Ait69, Ait83, BR97, Bar90, Bau85, BP79, BNS89, BP94, Bha07, Bow62a, Bow62b, Bro58, BS95, Cha93, Cul66, Cul72, Cul90, FJ11, Fie75, Fie86, Fin60, Fin66, Fin78, Fin11, Gan59a, Gan59b, Gan59c, Gan05, GLR83, GLR06, Gra69, Gra83, Hig08, Hou64, Hou75, Hou06, Lan69, LT85, Mac33, Mac46, Mac04, Min88, New72, Per52, Per91, PR71a, PR71b, Sen73b, Sen73a, Sen81, Sen06, Ser02, Sta69, ST68, Tan07a, TA32, Tur50, TA52, TA61, TA05, Wed34, Wed64, Wed05, WM67, GK60, Zur61, Zur64, ZF84, ZF86]. Matrix [BS70a, BS70b, Bar79, BL88, Bel70, Ber09, Bha87, Bha97, BR91, Don74, Fra68, Fra00, GLR82, GV96, Gra81, HJ85, HJ91, HJ12, Kel71, MM64, MM92, OCV11, Ort87, Rog80, Ste73, SS90, Var00, Zha02, Zha11, And87, Bar15]. Matrizen [Ait69, Zur50, Zur58, Zur61, Zur64, ZF84, ZF86]. mechanical [GK60]. mechanischer [GK60]. Melbourne [Bar15]. Methoden [CH24, CH37, CH42, CH68a, CH68b, ZF86]. Methods [Ano60, BS70a, BS70b, Bar79, Bar90, CH53, CH89, Cou89, FF63, Fad59, Tan07a, Tan07b, Tan07c, ZF86, CH24, CH37, CH42, CH68a, CH68b]. Models [Bap93]. Monotone [Don74]. Multilinear [Gre67, Gre78, Mar73, Mar75].

Negative [Sen73b, Sen73a, Sen81, Sen06]. Non [Sen73b, Sen73a, Sen81, Sen06]. Non-Negative [Sen73b, Sen73a, Sen81, Sen06]. Nonnegative [BR97, BP79, BNS89, BP94, FJ11, Min88]. Nonselfadjoint [GK69]. Normed [BD71]. Norms [BL88]. Notes [WM67, Fan59]. Numerical [BD71, BD73, Fie86, Hou64, Hou72, Hou75, Hou06, Tod62, YG72, YG73, YG88a, YG88b, ZF86]. Numerische [ZF86].

Operators [Bau85, BD71, GK69, Kat66, Kat76, Kat84, Kat95]. Order [Mui23, Mui60a]. Ordinary [Tan07b]. Oscillation [GK60]. Oszillationskerne [GK60]. Oszillationsmatrizen [GK60]. Ought [Gol04].

Part [ZF84, ZF86]. Partial [Cou89, Tan07c]. Perturbation [Bau85, Bha87, Kat66, Kat76, Kat84, Kat95, SS90]. Physicists [ZF84, ZF86]. Physics [CH24, CH37, CH42, CH53, CH68a, CH68b, CH89, Cou89]. Physik [CH24, CH37, CH42, CH68a, CH68b]. Physiker [ZF84, ZF86]. Polynomials [Bar83, GLR82]. Positive [Bha07]. Positivity [Kar68]. Presentation [GL06]. Press [Bar15]. Primer [Boa60, Boa72, Boa81, Boa96]. Principles [Rud53, Rud64, Rud76]. Problem [GL06, Hal67, Hal82, Par80, Par98, Wil65, Wil88]. Problems [LH74, LH95, OCV11, Zha09]. Products [GLR83, Gra81]. Properties [Fie75].

R [And87, Bar15]. Ranges [BD71, BD73]. Real [Boa60, Boa72, Boa81, Boa96]. Results [Zha11]. Review [And87, Bar15]. Right [Axl96]. Rings [Jac43]. Roger [And87, Bar15]. route [Kőn36].

Scalar [GLR83]. Schur [Zha05]. Schwingungen [GK60]. Sciences [Ano60, BP79, BP94]. Scientists [Bar79, Tan07a, Tan07b, Tan07c]. Second [Kap03, Ort87, Bar15]. Sets [Fan59, Lay82, Lay92, Lay07, Val64, Val76]. Sign [BS95]. Sign-Solvable [BS95]. Simultaneous [RR00]. small [GK60]. Social [Ano60]. Solvable [BS95]. Solving [LH74, LH95]. Some [Fie75]. Sound [Str78, Str96, Str45]. Space [Hal67, Hal82]. Spaces [BD71, Hal42, Hal58]. Special [Fie86]. Spectral [Fie75]. Squares [LH74, LH95]. Stability [BS70a, BS70b]. Statistics [Gra69, Gra83, PR71a, PR71b]. Streckenkomplexe [Kőn36]. Student [Gol04]. Students [Zha09]. Subspaces [GLR06]. Survey [MM64, MM92, Tod62, YG72, YG73, YG88a, YG88b]. Sylvester [Par06]. Symmetric [Par80, Par98]. Systematic [GL06]. Systeme [GK60]. Systems [Bar83, BNS89, BS95, HS74, GK60].

technical [Zur61, Zur64]. Techniques [Zha11]. technischen [Zur61, Zur64]. Teil [ZF84, ZF86]. Their [BL88, Fan59, Fie86, Lay82, Lay92, Lay07, Zur61, Zur64, ZF84, ZF86]. Theorie [Kőn36, Kőn50, Kőn54]. Theory [BS70a, BS70b, Bar75, BC85, Bau85, Ber09, Bro58, BR91, Fra68, Gan59a, Gan59b, Gan59c, Gan05, GK69, Hig08, Hou64, Hou75, Hou06, Jac43, Kat66, Kat76, Kat84, Kat95, Kel71, Kőn90, Lan69, LT85, Mac33, Mac46, Mac04, MM64, MM92, MOA11, Mui23, Mui60a, Mui60b, Ort87, Per52, Per91, Sen73b, Sen73a, Ser02, SS90, Str78, Str96, Str45, TA32, Tur50, TA52, TA61, TA05, WM67, Zha11, Fra00, Kőn36, Kőn50, Kőn54, MO79]. Topics [HJ91, Kel71, OCV11, WM67]. Topologie [Kőn36]. topology [Kőn36]. Total [Kar68]. Totally [FJ11]. Transformations [CM79, CM91, CM09, Cul66, Cul72, Cul90, Fin60, Fin66, Fin78, Fin11, Sta69]. Transforms [Tan07b]. Treatise [Mui60b]. Triangularization [RR00].

Undergraduate [C^{+}02]. unendlichen [Kőn36, Kőn50, Kőn54]. University [Bar15]. Uses [ZF84, ZF86].

Variational [Tan07c]. Vector [Hal42, Hal58, Tan07b]. vibrations [GK60]. Victorian [Cri06, Par06].

Weaving [OCV11]. werke [HS96]. Weyr [OCV11]. Wielandt [HS96]. Works [HS96]. World [Par06].

References

Aitken:1939:DM

[Ait39] A. C. (Alexander Craig) Aitken. Determinants and Matrices. University mathematical texts. Oliver and Boyd, Edinburgh, UK, 1939. vii + 135 pp. LCCN QA191 .A5.

Aitken:1946:DM

[Ait46] A. C. (Alexander Craig) Aitken. Determinants and Matrices. Oliver and Boyd, Edinburgh and London, UK, fourth edition, 1946. vii + 143 + 1 pp. LCCN QA191 .A5 1946.

Aitken:1948:DM

[Ait48] A. C. (Alexander Craig) Aitken. Determinants and Matrices. University mathematical texts. Oliver and Boyd, Edinburgh, UK, fifth edition, 1948. vii + 143 pp. LCCN QA191 .A5 1948.

Aitken:1954:DM

[Ait54] A. C. (Alexander Craig) Aitken. Determinants and Matrices. University mathematical texts. Oliver and Boyd, Edinburgh, UK, eighth edition, 1954. 144 pp. LCCN QA191 .A5 1954.

Aitken:1956:DM

[Ait56] A. C. (Alexander Craig) Aitken. Determinants and Matrices. University mathematical texts. Oliver and Boyd, Edinburgh, UK, ninth edition, 1956. 144 pp. LCCN QA191 .A5 1956.

Aitken:1969:DMG

[Ait69] A. C. (Alexander Craig) Aitken. Determinanten und Matrizen. (German) [Determinants and Matrices], volume 293 of B.I.- Hochschultaschenbücher. Bibliographisches Institut, Mannheim, Germany, 1969. 142 pp. LCCN QA191 .A515. Translation to German by Winfried Nilson of [Ait56].

Aitken:1983:DM

[Ait83] A. C. (Alexander Craig) Aitken. Determinants and Matrices. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, revised edition, 1983. ISBN 0-313-23294-6. vii + 144 pp. LCCN QA191 .A5 1983.

Ando:1987:RBA

[And87] T. Ando. Review of Matrix Analysis, by Roger A. Horn and Charles R. Johnson. Linear Algebra and its Applications, 90(??):245–247, May 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0024379587903181.

Anonymous:1960:MMS

[Ano60] Anonymous. Mathematical Methods in the Social Sciences, 1959. Proceedings of the first Stanford Symposium. Stanford Mathematical Studies in the Social Sciences, IV. Stanford University Press, Stanford, CA, USA, 1960. viii + 365 pp. LCCN ????

Axler:1996:LAD

[Axl96] Sheldon Axler. Linear Algebra Done Right. Undergraduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94595-4 (hardcover), 0-387-94596-2 (paperback). xviii + 238 pp. LCCN QA184 .A96

1996. URL http://www.zentralblattmath.org/zmath/en/search/
?an=0843.15002.

Bapat:1993:LAL

[Bap93] R. B. Bapat. Linear Algebra and Linear Models. Texts and Readings in Mathematics. Hindustan Book Agency, Delhi, India, 1993. ISBN 81-85931-00-3. viii + 124 pp. LCCN QA184 .B37 1993. URL http:/

/www.zentralblattmath.org/zmath/en/search/?an=0834.62062.

Barnett:1975:IMC

[Bar75] Stephen Barnett. Introduction to Mathematical Control Theory. Oxford Applied Mathematics and Computing Science Series. Clarendon Press, Oxford, UK and New York, NY, USA, 1975. ISBN 0-19-859618-9. viii + 264 pp. LCCN QA402.3 .B347.

Barnett:1979:MME

[Bar79] S. Barnett. Matrix Methods for Engineers and Scientists. McGraw-Hill, New York, NY, USA, 1979. ISBN 0-07-084084-9. xi + 185 pp. LCCN QA188 .B37.

Barnett:1983:PLC

[Bar83] Stephen Barnett. Polynomials and Linear Control Systems, volume 77 of Monographs and Textbooks in Pure and Applied Mathematics. Marcel Dekker, Inc., New York, NY, USA, 1983. ISBN 0-8247-1898-4 (hardcover), 0-8247-2898-4. xi + 452 pp. LCCN QA161.P59 B37 1983.

Barnett:1990:MMA

[Bar90] Stephen Barnett. Matrices: Methods and Applications. Oxford Applied Mathematics and Computing Science Series. Clarendon Press, Oxford, UK and New York, NY, USA, 1990. ISBN 0-19-859665-0, 0-19-859680-4 (paperback). xvi + 450 pp. LCCN QA188 .B36

1990. URL http://catdir.loc.gov/catdir/enhancements/fy0603/
89023942-d.html;   http://catdir.loc.gov/catdir/
enhancements/fy0603/89023942-t.html;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0706.15001.

Bart:2015:RBA

[Bar15] Harm Bart. Review of Matrix Analysis, second edition, Roger A. Horn, Charles R. Johnson. Cambridge University Press, Cambridge, Melbourne etc. (2013), ISBN: 978-0-521-54823-6. Linear Algebra and its Applications, 466(??):527–529, February 1, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379514006922.

Baumgartel:1985:APT

[Bau85] H. Baumgärtel. Analytic Perturbation Theory for Matrices and Operators, volume 15 of Operator Theory: Advances and Applications. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1985. ISBN 3-7643-1664-0. 427 pp. LCCN QA188 .B38 1985.

URL http://www.gbv.de/dms/hbz/toc/ht002465460.pdf.

Beckenbach:1965:I

[BB65] Edwin F. Beckenbach and Richard Bellman. Inequalities, volume 30 of Second revised printing. Ergebnisse der Mathematik und ihrer Grenzgebiete. Neue Folge. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1965. xi + 198 pp. LCCN ????

Barnett:1985:IMC

[BC85] Stephen Barnett and R. G. Cameron. Introduction to Mathematical Control Theory. Oxford applied mathematics and computing science series. Clarendon Press, Oxford, UK and New York, NY, USA, second edition, 1985. ISBN 0-19-859640-5, 0-19-859639-1 (paperback). xi + 404 pp. LCCN QA402.3 .B347 1985. URL http://www.gbv.de/

dms/hbz/toc/ht002139017.pdf;   http://www.loc.gov/catdir/
enhancements/fy0640/85021589-d.html.

Bonsall:1971:NRO

[BD71] F. F. Bonsall and J. Duncan. Numerical Ranges of Operators on Normed Spaces and of Elements of Normed Algebras, volume 2 of London Mathematical Society Lecture Note Series. Cambridge University Press, Cambridge, UK, 1971. ISBN 0-521-07988-8 (paperback), 1-107-35989-9 (e-book). iv + 142 pp. LCCN QA322

.B65. URL http://www.loc.gov/catdir/enhancements/fy0808/
71128498-d.html;   http://www.loc.gov/catdir/enhancements/
fy0808/71128498-t.html.

Bonsall:1973:NRI

[BD73] F. F. Bonsall and J. Duncan. Numerical Ranges. II, volume 10 of London Mathematical Society Lecture Notes. Cambridge University Press, Cambridge, UK, 1973. ISBN 0-521-20227-2 (paperback), 0-511-66251-3 (e-book). vii + 179 pp. LCCN QA322.2 .B66. URL http://www.loc.

gov/catdir/enhancements/fy0827/73179782-d.html;   http://
www.loc.gov/catdir/enhancements/fy0827/73179782-t.html.

Bellman:1970:IMA

[Bel70] Richard Bellman. Introduction to Matrix Analysis. McGraw-Hill, New York, NY, USA, second edition, 1970. xxiii + 403 pp. LCCN ????

Bernstein:2009:MMT

[Ber09] Dennis S. Bernstein. Matrix Mathematics: Theory, Facts, and Formulas. Princeton University Press, Princeton, NJ, USA, second edition, 2009. ISBN 0-691-13287-9 (hardcover), 0-691-14039-1 (paperback), 1-4008-3334-5 (e-book). xlii + 1139 pp. LCCN QA188 .B475 2009;

QA188  .B475X  2009

(LC). URL http://www.jstor.org/stable/10.2307/j.ctt7t833;
 http://www.loc.gov/catdir/toc/ecip0826/2008036257.html.

Bhatia:1987:PBM

[Bha87] Rajendra Bhatia. Perturbation Bounds for Matrix Eigenvalues, volume 162 of Pitman Research Notes in Mathematics Series. Longman Scientific and Technical, Harlow, Essex, UK, 1987. ISBN 0-470-20917-8 (paperback), 0-582-01379-8 (paperback). ISSN 0269-3674. viii +

129 pp. LCCN QA193. URL http://www.gbv.de/dms/hbz/toc/
ht002967626.pdf;   http://www.zentralblatt-math.org/zmath/
en/search/?an=0696.15013.

Bhatia:1997:MA

[Bha97] Rajendra Bhatia. Matrix Analysis, volume 169 of Graduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94846-5 (print), 1-4612-0653-7 (e-book). xi + 347 pp. LCCN QA188 .B485 1996. URL http://

catdir.loc.gov/catdir/enhancements/fy0818/96032217-d.html;
 http://www.gbv.de/dms/goettingen/213798093.pdf.

Bhatia:2007:PDM

[Bha07] Rajendra Bhatia. Positive Definite Matrices. Princeton Series in Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12918-5. x + 254 pp. LCCN QA188 .B488

2007. URL http://www.loc.gov/catdir/enhancements/fy0713/
2006050375-d.html;   http://www.loc.gov/catdir/
enhancements/fy0713/2006050375-t.html;   http://www.loc.
gov/catdir/enhancements/fy0734/2006050375-b.html.

Belitskii:1988:MNT

[BL88] Genrikh Ruvimovich Belitskiĭ and Yuri Ilich Lyubich. Matrix Norms and Their Applications, volume 36 of Operator Theory: Advances and Applications. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1988. ISBN 3-7643-2220-9 (Basel), 0-8176-2220-9 (Boston). viii + 209 pp. LCCN QA188 .B4313

1988. URL http://www.gbv.de/dms/hbz/toc/ht003195543.pdf;
http://www.zentralblatt-
math.org/zmath/en/search/?an=0645.15019. Translated from the

Russian by A. Iacob.

Berman:1989:NMD

[BNS89] Abraham Berman, Michael Neumann, and Ronald J. Stern. Nonnegative Matrices in Dynamic Systems. Pure and Applied Mathematics (New York). Wiley, New York, NY, USA, 1989. ISBN 0-471-62074-2 (hardcover). xxii + 167 pp. LCCN QA188

.B466 1989. URL http://catdir.loc.gov/catdir/description/
wiley031/88033934.html;   http://www.gbv.de/dms/hbz/toc/
ht003360524.pdf;   http://www.loc.gov/catdir/description/
wiley031/88033934.html;   http://www.loc.gov/catdir/toc/
onix02/88033934.html;   https://www.math.utah.edu/pub/
bibnet/subjects/matrix-analysis-2ed.bib;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0723.93013.

Boas:1960:PRF

[Boa60] Ralph P. (Ralph Philip) Boas. A Primer of Real Functions, volume 13 of The Carus mathematical monographs. Mathematical Association of America, Washington, DC, USA, 1960. xiii + 189 pp. LCCN QA331 .B646.

Boas:1972:PRF

[Boa72] Ralph P. (Ralph Philip) Boas. A Primer of Real Functions, volume 13 of The Carus mathematical monographs. Mathematical Association of America, Washington, DC, USA, second edition, 1972. ISBN 0-88385-013-3. xi + 196 pp. LCCN QA331.5 .B57 1972.

Boas:1981:PRF

[Boa81] Ralph P. (Ralph Philip) Boas. A Primer of Real Functions, volume 13 of The Carus mathematical monographs. Mathematical Association of America, Washington, DC, USA, third edition, 1981. ISBN 0-88385-022-2. ISSN 0069-0813. xi + 232 pp. LCCN QA331.5 .B57 1981. URL

http://www.gbv.de/dms/bowker/toc/9780883850220.pdf.

Boas:1996:PRF

[Boa96] Ralph P. (Ralph Philip) Boas. A Primer of Real Functions, volume 13 of Carus Mathematical Monographs. Mathematical Association of America, Washington, DC, USA, fourth edition, 1996. ISBN 0-88385-029-X. xiv + 305 pp. LCCN QA331.5 .B57 1996. URL http:

//catdir.loc.gov/catdir/description/cam028/96077785.html;
 http://catdir.loc.gov/catdir/toc/cam027/96077785.html. Re-

vised and with a preface by Harold P. Boas.

Bowman:1962:IDMa

[Bow62a] Frank Bowman. An Introduction to Determinants and Matrices: Determinants and Matrices. Applied mathematics series. Van Nostrand, Princeton, NJ, USA, 1962. 163 pp. LCCN QA191.

Bowman:1962:IDMb

[Bow62b] Frank Bowman. An Introduction to Determinants and Matrices: Determinants and Matrices. Applied mathematics series. English Universities Press, London, UK, 1962. ix + 163 pp. LCCN QA191 .B6.

Berman:1979:NMM

[BP79] Abraham Berman and Robert J. Plemmons. Nonnegative Matrices in the Mathematical Sciences. Computer Science and Applied Mathematics. Academic Press, New York, NY, USA, 1979. ISBN 0-12-092250-9. xviii + 316 pp. LCCN QA188.B47 1979. URL http://www.

sciencedirect.com/science/book/9780120922505.

Berman:1994:NMM

[BP94] Abraham Berman and Robert J. Plemmons. Nonnegative Matrices in the Mathematical Sciences, volume 9 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-321-8. xx + 340 pp. LCCN QA188 .B47

1994. URL http://www.loc.gov/catdir/enhancements/fy0708/
94037449-d.html;   http://www.
loc.gov/catdir/enhancements/fy0708/94037449-t.html. Revised

reprint of [BP79].

Brualdi:1991:CMT

[BR91] Richard A. Brualdi and Herbert J. Ryser. Combinatorial Matrix Theory, volume 39 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-32265-0. x + 367 pp. LCCN QA188 .B78 1991. URL http://www.loc.gov/

catdir/description/cam024/90020210.html;
http://www.loc.gov/catdir/toc/cam024/90020210.html.

Bapat:1997:NMA

[BR97] R. B. Bapat and T. E. S. Raghavan. Nonnegative Matrices and Applications, volume 64 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-57167-7 (hardcover), 1-107-09513-1, 0-511-52997-X (e-book). xiv + 336 pp. LCCN QA188 .B355 1997.

Browne:1958:ITD

[Bro58] Edward Tankard Browne. Introduction to the Theory of Determinants and Matrices. University of North Carolina Press, Chapel Hill, NC, USA, 1958. 270 pp. LCCN QA263 .B7.

Barnett:1970:MMSa

[BS70a] S. Barnett and C. Storey. Matrix Methods in Stability Theory. Applications of mathematics series. Barnes and Noble, New York, NY, USA, 1970. viii + 148 pp. LCCN QA263 .B33 1970.

Barnett:1970:MMSb

[BS70b] S. Barnett and C. Storey. Matrix Methods in Stability Theory. Applications of mathematics series. Nelson, London, UK, 1970. ISBN 0-17-761617-2. viii + 148 pp. LCCN QA263 .B33 1970.

Brualdi:1995:MSS

[BS95] Richard A. Brualdi and Bryan L. Shader. Matrices of Sign-Solvable Linear Systems, volume 116 of Cambridge Tracts in Mathematics. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-48296-8. xii + 298 pp. LCCN QA188 .B79 1995. URL http://www.loc.

gov/catdir/description/cam026/94040931.html;
http://www.loc.gov/catdir/toc/cam024/94040931.html.

Carlson:2002:LAG

[C^{+}02] David (David H.) Carlson et al., editors. Linear Algebra Gems: Assets for Undergraduate Mathematics, volume 59 of MAA notes. Mathematical Association of America, Washington, DC, USA, 2002. ISBN 0-88385-170-9 (paperback). xvi + 328 pp. LCCN QA184.2 .L55 2002.

Courant:1924:MMP

[CH24] Richard Courant and David Hilbert. Methoden der mathematischen Physik. (German) [Methods of Mathematical Physics], volume 1 of Die Grundlehren der mathematischen Wissenschaften (12). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1924. xiii + 450 pp.

Courant:1937:MMP

[CH37] Richard Courant and David Hilbert. Methoden der mathematischen Physik. (German) [Methods of Mathematical Physics], volume 2 of Die Grundlehren der mathematischen Wissenschaften (48). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1937. xvi + 549 pp.

Courant:1942:MMP

[CH42] Richard Courant and David Hilbert. Methoden der mathematischen Physik. (German) [Methods of Mathematical Physics], volume 2 of Die Grundlehren der mathematischen Wissenschaften (48). Interscience, New York, NY, USA, 1942. xiv + 549 pp. LCCN ????

Courant:1953:MMP

[CH53] Richard Courant and David Hilbert. Methods of Mathematical Physics, volume I. Interscience Publishers, Inc., New York, NY, USA, 1953. ISBN 0-471-50447-5, 0-585-29428-3 (e-book). xv + 561 pp. LCCN ????

Courant:1968:MMPa

[CH68a] Richard Courant and David Hilbert. Methoden der mathematischen Physik. (German) [Methods of Mathematical Physics], volume 1 of Heidelberger Taschenbücher (30). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1968. ISBN 0-387-04177-X (New York), 3-540-04177-X (Berlin). xv + 469 pp. LCCN ????

Courant:1968:MMPb

[CH68b] Richard Courant and David Hilbert. Methoden der mathematischen Physik. (German) [Methods of Mathematical Physics], volume 2 of Heidelberger Taschenbücher (31). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1968. ISBN 0-387-04178-8, 3-540-04178-8. xvi + 549 pp. LCCN ????

Courant:1989:MMPa

[CH89] Richard Courant and David Hilbert. Methods of Mathematical Physics, volume 1 of Wiley classics library. Wiley, New York, NY, USA, 1989. ISBN 0-471-50447-5, 3-527-61721-3 (e-book). xv + 560 pp.

LCCN QA401 .C72413 1989. URL http://onlinelibrary.wiley.
com/book/10.1002/9783527617210;   http://www.gbv.de/dms/
bowker/toc/9780471504474.pdf;   http://www.zentralblatt-
math.org/zmath/en/search/?an=0729.00007.

Chatelin:1993:EM

[Cha93] Françoise Chatelin. Eigenvalues of Matrices. Pure and applied mathematics. Wiley, New York, NY, USA, 1993. ISBN 0-471-93538-7. xviii + 382 pp. LCCN QA188 .C44 1993. URL http://www.loc.gov/

catdir/enhancements/fy0707/93003430-d.html;   http:/
/www.loc.gov/catdir/toc/onix04/93003430.html. With exercises

by Mario Ahués and the author, Translated from the French and with additional material by Walter Ledermann.

Childs:1979:CIH

[Chi79] Lindsay N. Childs. A Concrete Introduction to Higher Algebra. Undergraduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-90333-X (New York), 3-540-90333-X (Berlin). xiv + 338 pp. LCCN QA155 .C53.

Childs:1995:CIH

[Chi95] Lindsay N. Childs. A Concrete Introduction to Higher Algebra. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1995. ISBN 0-387-94484-2 (hardcover), 1-4419-8702-9 (e-book). xv + 522 pp. LCCN QA155 .C53 1995.

Childs:2009:CIH

[Chi09] Lindsay N. Childs. A Concrete Introduction to Higher Algebra. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2009. ISBN 0-387-74527-0, 0-387-74725-7 (e-book). xiv + 603 pp. LCCN QA155 .C53 2009.

Campbell:1979:GIL

[CM79] S. L. (Stephen La Vern) Campbell and C. D. (Carl Dean) Meyer, Jr. Generalized Inverses of Linear Transformations. Surveys and reference works in mathematics. Pitman Publishing Ltd., London, UK, 1979. ISBN 0-273-08422-4. xi + 272 pp. LCCN QA188 .C36. Reprinted in [CM91].

Campbell:1991:GIL

[CM91] S. L. (Stephen La Vern) Campbell and C. D. (Carl Dean) Meyer, Jr. Generalized Inverses of Linear Transformations. Dover, New York, NY, USA, 1991. ISBN 0-486-66693-X (paperback). xi + 272 pp. LCCN

QA188 .C36 1991. URL http://www.loc.gov/catdir/description/
dover032/90023852.html.

Campbell:2009:GIL

[CM09] S. L. (Stephen La Vern) Campbell and C. D. (Carl Dean) Meyer, Jr. Generalized Inverses of Linear Transformations, volume 56 of Classics in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2009. ISBN 0-89871-671-3, 0-89871-904-6 (e-book). xx + 272 pp. LCCN QA188 .C36

2009. URL http://epubs.siam.org/ebooks/siam/classics_in_
applied_mathematics/cl56;   http://www.loc.gov/catdir/
enhancements/fy0916/2008046428-b.html;   http://www.loc.
gov/catdir/enhancements/fy0916/2008046428-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0916/2008046428-t.html.

Conway:1990:CFA

[Con90] John B. Conway. A Course in Functional Analysis, volume 96 of Graduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1990. ISBN 0-387-97245-5 (New York), 3-540-97245-5 (Berlin). xvi + 399 pp. LCCN QA320 .C658

1990. URL http://catdir.loc.gov/catdir/enhancements/fy0813/
97122669d.html;   http://catdir.loc.gov/catdir/
enhancements/fy0813/97122669t.html;   http://www.gbv.de/
dms/bowker/toc/9780387972459.pdf;   http://www.gbv.de/dms/

ilmenau/toc/025248731.PDF;   http://www.loc.gov/catdir/
enhancements/fy0813/97122669-d.html;   http://www.loc.gov/
catdir/enhancements/fy0813/97122669-t.html;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0706.46003.

Courant:1989:MMPb

[Cou89] Richard Courant. Methods of Mathematical Physics: Partial Differential Equations, volume 2 of Wiley classics library. Interscience Publishers, New York, NY, USA, 1989. ISBN 0-471-50439-4 (vol. 2). xxii + 830 pp. LCCN QA401 .C72413 1989. URL http://onlinelibrary.

wiley.com/book/10.1002/9783527617234;   http://www.loc.
gov/catdir/enhancements/fy0607/89112829-b.html;   http://
www.loc.gov/catdir/enhancements/fy0607/89112829-d.html;
http://www.loc.gov/catdir/toc/onix05/89112829.html.

Crilly:2006:ACM

[Cri06] A. J. (Tony) Crilly. Arthur Cayley: Mathematician Laureate of the Victorian Age. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8011-4. xxiv + 610 pp. LCCN QA29 .C39 C75

2006. URL http://www.loc.gov/catdir/bios/jhu052/2004005682.
html;   http://www.loc.gov/catdir/description/jhu051/
2004005682.html;
http://www.loc.gov/catdir/toc/fy0612/2004005682.html.

Cullen:1966:MLT

[Cul66] Charles G. Cullen. Matrices and Linear Transformations. Addison-Wesley, Reading, MA, USA, 1966. viii + 227 pp. LCCN QA263 .C8.

Cullen:1972:MLT

[Cul72] Charles G. Cullen. Matrices and Linear Transformations. Addison-Wesley series in mathematics. Addison-Wesley, Reading, MA, USA, second edition, 1972. ISBN 0-201-01209-X. xii + 318 pp. LCCN QA263 .C8 1972.

Cullen:1990:MLT

[Cul90] Charles G. Cullen. Matrices and Linear Transformations. Dover, New York, NY, USA, second edition, 1990. ISBN 0-486-66328-0. xii + 318

pp. LCCN QA188 .C85 1990. URL http://www.loc.gov/catdir/
description/dover032/89025677.html;   http://www.loc.gov/
catdir/enhancements/fy1318/89025677-t.html.

Donoghue:1974:MMF

[Don74] William F. (William Francis) Donoghue, Jr. Monotone Matrix Functions and Analytic Continuation, volume 207 of Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebeite. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 0-387-06543-1. v + 182 pp. LCCN QA331 .D687.

Faddeeva:1959:CML

[Fad59] V. N. Faddeeva. Computational Methods of Linear Algebra. Dover, New York, NY, USA, 1959. xi + 252 pp. LCCN QA251 .F313.

Fan:1959:CST

[Fan59] Ky Fan. Convex Sets and Their Applications [lecture notes]. Applied Mathematics Division, Argonne National Laboratory, Argonne, IL, USA, 1959. ???? pp. LCCN ????

Faddeev:1963:CML

[FF63] D. K. (Dmitriĭ Konstantinovich) Faddeev and V. N. Faddeeva. Computational Methods of Linear Algebra. A Series of undergraduate books in mathematics. W. H. Freeman, New York, NY, USA, 1963. xi + 621 pp. LCCN QA251 .F283.

Fiedler:1975:SPS

[Fie75] Miroslav Fiedler. Spectral properties of some classes of matrices. Lecture Notes 75.01R, Chalmers University of Technology and the University of Göteborg, Göteborg, Sweden, 1975. ???? pp.

Fiedler:1986:SMT

[Fie86] Miroslav Fiedler. Special Matrices and Their Applications in Numerical Mathematics. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1986. ISBN 90-247-2957-2. xi + 307 pp. LCCN QA188 .F5313

1986. URL http://www.gbv.de/dms/hbz/toc/ht002867827.pdf;
http://www.zentralblatt-math.org/zmath/en/search/?an=0677.

65019. Translated from the Czech Speciální matice a jejich použití v numerické matematice by Petr Přikryl and Karel Segeth.

Finkbeiner:1960:IML

[Fin60] Daniel T. (Daniel Talbot) Finkbeiner II. Introduction to Matrices and Linear Transformations. A Series of undergraduate books in mathematics. W. H. Freeman, New York, NY, USA, 1960. 248 pp. LCCN QA263 .F45.

Finkbeiner:1966:IML

[Fin66] Daniel T. (Daniel Talbot) Finkbeiner II. Introduction to Matrices and Linear Transformations. A Series of books in mathematics. W. H. Freeman, New York, NY, USA, second edition, 1966. xi + 297 pp. LCCN QA263 .F45 1966.

Finkbeiner:1978:IML

[Fin78] Daniel T. (Daniel Talbot) Finkbeiner II. Introduction to Matrices and Linear Transformations. A Series of books in the mathematical sciences. W. H. Freeman, New York, NY, USA, third edition, 1978. ISBN 0-7167-0084-0. xii + 462 pp. LCCN QA188 .F56 1978.

Finkbeiner:2011:IML

[Fin11] Daniel T. (Daniel Talbot) Finkbeiner II. Introduction to Matrices and Linear Transformations. Dover, New York, NY, USA, dover edition, 2011. ISBN 0-486-48159-X. xii + 462 pp. LCCN QA188 .F56

2011. URL http://www.loc.gov/catdir/enhancements/fy1106/
2010048976-d.html;   http://www.loc.gov/catdir/
enhancements/fy1318/2010048976-t.html.

Fallat:2011:TNM

[FJ11] Shaun M. Fallat and Charles R. Johnson. Totally Nonnegative Matrices. Princeton Series in Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-12157-5 (hardcover), 1-4008-3901-7 (e-book). xv + 248 pp. LCCN QA188 .F35 2011. URL

http://www.jstor.org/stable/10.2307/j.ctt7scff.

Franklin:1968:MT

[Fra68] Joel N. Franklin. Matrix Theory. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1968. xii + 292 pp. LCCN QA263 .F72.

Franklin:2000:MT

[Fra00] Joel N. Franklin. Matrix theory. Dover, New York, NY, USA, 2000. ISBN 0-486-41179-6 (paperback). xii + 292 pp. LCCN QA188 .F66

2000. URL http://www.loc.gov/catdir/description/dover031/
99058316.html;
http://www.loc.gov/catdir/toc/dover031/99058316.html.

Gantmacher:1959:ATM

[Gan59a] F. R. (Feliks Ruvimovich) Gantmacher. Applications of the Theory of Matrices. Translated by J. L. Brenner, with the assistance of D. W.

Bushaw and S. Evanusa. Interscience Publishers, Inc., New York, NY, USA, 1959. ix + 317 pp. LCCN QA263 .G3532.

Gantmacher:1959:TMa

[Gan59b] F. R. (Feliks Ruvimovich) Gantmacher. The Theory of Matrices, volume 1. Chelsea Publishing Company, New York, NY, USA, 1959. x + 374 pp. LCCN QA263.G35 1960 vol. 1.

Gantmacher:1959:TMb

[Gan59c] F. R. (Feliks Ruvimovich) Gantmacher. The Theory of Matrices, volume 2. Chelsea Publishing Company, New York, NY, USA, 1959. ix + 276 pp. LCCN QA263.G35 1960 vol. 2.

Gantmakher:2005:ATM

[Gan05] F. R. (Feliks Ruvimovich) Gantmakher. Applications of the Theory of Matrices. Dover, New York, NY, USA, 2005. ISBN 0-486-44554-2 (paperback). ix + 317 pp. LCCN QA263 .G3532

2005. URL http://www.loc.gov/catdir/enhancements/fy0624/

2005047409-d.html. Translated by J. L. Brenner and with the assistance of D. W. Busaw and S. Evanusa.

Gantmacher:1960:OOK

[GK60] F. R. (Feliks Ruvimovich) Gantmacher and M. G. Kreĭn. Oszillationsmatrizen, Oszillationskerne und kleine Schwingungen mechanischer Systeme. (German) [Oscillation matrices, oscillation kernels, and small vibrations of mechanical systems], volume 5 of Wissenschaftliche Bearbeitung der deutschen Ausgabe: Alfred Stöhr. Mathematische Lehrbücher und Monographien, I. Abteilung, Bd. V. Akademie-Verlag, Berlin, Germany, 1960. x + 359 pp. LCCN QA871 .G315.

Gohberg:1969:ITL

[GK69] I. C. Gohberg and M. G. Kreĭn. Introduction to the Theory of Linear Nonselfadjoint Operators, volume 18 of Translations of Mathematical Monographs. American Mathematical Society, Providence, RI, USA, 1969. ISBN 0-8218-1568-7. xv + 378 pp. LCCN QA320 .G6313. Translated from the Russian by A. Feinstein.

Glazman:2006:FDL

[GL06] I. M. Glazman and Ju. I. Ljubič. Finite-Dimensional Linear Analysis: a Systematic Presentation in Problem Form. Dover, New York, NY, USA, 2006. ISBN 0-486-45332-4. xx + 520 pp. LCCN QA320

.G5513 2006. URL http://www.loc.gov/catdir/enhancements/
fy0643/2006045447-d.html;   http://www.

zentralblattmath.org/zmath/en/search/?an=1115.46001. Trans-

lated from the Russian Konechnomernyi lineinyi analiz v zadachakh, and edited by G. P. Barker and G. Kuerti. Reprint of the 1974 MIT Press (Cambridge, MA, USA) edition.

Gohberg:1982:MP

[GLR82] I. Gohberg, P. Lancaster, and L. Rodman. Matrix Polynomials. Computer Science and Applied Mathematics. Academic Press, New York, NY, USA, 1982. ISBN 0-12-287160-X. xiv + 409 pp. LCCN

QA188 .G64 1982. URL http://www.loc.gov/catdir/description/
els033/81022792.html;
http://www.loc.gov/catdir/toc/els031/81022792.html.

Gohberg:1983:MIS

[GLR83] I. Gohberg, P. Lancaster, and L. Rodman. Matrices and Indefinite Scalar Products, volume 8 of Operator Theory: Advances and Applications. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1983. ISBN 3-7643-1527-X. xvii + 374 pp. LCCN QA329.2

.G63 1983. URL http://www.gbv.de/dms/ilmenau/toc/024274224.
PDF.

Gohberg:2006:ISM

[GLR06] Israel Gohberg, Peter Lancaster, and Leiba Rodman. Invariant Subspaces of Matrices with Applications, volume 51 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2006. ISBN 0-89871-608-X. xxi + 692

pp. LCCN QA322 .G649 2006. URL http://www.loc.gov/catdir/
enhancements/fy0665/2006042260-d.html;   http://www.loc.
gov/catdir/enhancements/fy0665/2006042260-t.html. Reprint of

the 1986 original.

Golan:2004:LAB

[Gol04] Jonathan S. Golan. The Linear Algebra a Beginning Graduate Student Ought to Know, volume 27 of Kluwer Texts in the Mathematical Sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2004. ISBN 1-4020-1824-X (hardcover). x + 406 pp. LCCN QA184.2 .G65 2004.

Graybill:1969:IMA

[Gra69] Franklin A. (Franklin Arno) Graybill. Introduction to Matrices with Applications in Statistics. Wadsworth International Group, Belmont, CA, USA, 1969. 372 pp. LCCN QA263 .G67.

Graham:1981:KPM

[Gra81] Alexander Graham. Kronecker Products and Matrix Calculus: with Applications. Ellis Horwood Series in Mathematics and its Applications. Ellis Horwood, New York, NY, USA, 1981. ISBN 0-85312-391-8 (library), 0-85312-427-2 (student), 0-470-27300-3 (Halsted Press). 130 pp. LCCN QA188 .G698 1981.

Graybill:1983:MAS

[Gra83] Franklin A. (Franklin Arno) Graybill. Matrices with Applications in Statistics. Wadsworth statistics/probability series. Wadsworth International Group, Belmont, CA, USA, second edition, 1983. ISBN 0-534-98038-4. xi + 461 pp. LCCN QA188 .G715 1983.

Greub:1967:MA

[Gre67] Werner Hildbert Greub. Multilinear Algebra, volume 136 of Die Grundlehren der Mathematischen Wissenschaften. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. x + 224 pp. LCCN QA251 .G72.

Greub:1978:MA

[Gre78] Werner Hildbert Greub. Multilinear Algebra. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1978. ISBN 0-387-90284-8. vii + 294 pp. LCCN QA184 .G74 1978.

Golub:1996:MC

[GV96] Gene H. Golub and Charles F. Van Loan. Matrix Computations. Johns Hopkins Studies in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, third edition, 1996. ISBN 0-8018-5413-X (hardcover), 0-8018-5414-8 (paperback). xxvii + 694 pp.

LCCN QA188 .G65 1996. URL http://www.loc.gov/catdir/bios/
jhu052/96014291.html;   http://www.loc.gov/catdir/
description/jhu051/96014291.html.

Halmos:1942:FDV

[Hal42] Paul R. (Paul Richard) Halmos. Finite-Dimensional Vector Spaces. The University Series in Undergraduate Mathematics. Princeton University Press, Princeton, NJ, USA, 1942. v + 196 pp. LCCN QA601 .H23 1942.

Halmos:1958:FDV

[Hal58] Paul R. (Paul Richard) Halmos. Finite-Dimensional Vector Spaces. The University Series in Undergraduate Mathematics. D. Van Nostrand Co., Inc., Princeton / Toronto / New York / London, second edition, 1958. viii + 200 pp. LCCN QA261 .H33 1958.

Halmos:1967:HSP

[Hal67] Paul R. (Paul Richard) Halmos. A Hilbert Space Problem Book. D. Van Nostrand Co., Inc., Princeton / Toronto / New York / London, 1967. xvii + 365 pp. LCCN QA322 .H3.

Halmos:1982:HSP

[Hal82] Paul R. (Paul Richard) Halmos. A Hilbert Space Problem Booko, volume 19 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1982. ISBN 0-387-90685-1. xvii + 369 pp. LCCN QA322.4 .H34

1982. URL http://www.loc.gov/catdir/enhancements/fy0905/
82000763-d.html;   http://www.loc.gov/catdir/enhancements/
fy0905/82000763-t.html.

Higham:2008:FMT

[Hig08] Nicholas J. Higham. Functions of Matrices: Theory and Computation. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2008. ISBN 0-89871-646-2 (paperback). xx + 425

pp. LCCN QA188 .H53 2008. URL http://www.loc.gov/catdir/
enhancements/fy0834/2007061811-b.html;   http://www.
loc.gov/catdir/enhancements/fy0834/2007061811-d.html. The-

ory and computation.

Horn:1985:MA

[HJ85] Roger A. Horn and Charles R. Johnson. Matrix Analysis. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30586-1 (hardcover), 0-521-38632-2 (paperback), 0-511-81081-

4 (e-book).  xiii + 561 pp.  LCCN QA188 .H66 1985.

URL http://www.gbv.de/dms/bowker/toc/9780521305860.pdf;
http://www.gbv.de/dms/hbz/toc/ht002524989.pdf;
http://www.gbv.de/dms/ilmenau/toc/196021510.PDF;   http://
www.gbv.de/dms/ilmenau/toc/576799351.PDF;   http://www.
loc.gov/catdir/description/cam023/85007736.html;   http://
www.loc.gov/catdir/toc/cam023/85007736.html;   http://www.
zentralblattmath.org/zmath/en/search/?an=0704.15002.

Horn:1991:TMA

[HJ91] Roger A. Horn and Charles R. Johnson. Topics in Matrix Analysis. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-30587-X. viii + 607 pp. LCCN QA188 .H664 1991. URL http://www.loc.

gov/catdir/description/cam023/86023310.html;
http://www.loc.gov/catdir/toc/cam023/86023310.html.

Horn:2012:MA

[HJ12] Roger A. Horn and Charles R. (Charles Royal) Johnson. Matrix Analysis. Cambridge University Press, Cambridge, UK, second edition, 2012. ISBN 0-521-83940-8 (hardcover), 0-521-54823-3 (paperback), 1-283-74139-3, 1-139-77904-4, 1-139-77600-2 (e-book), 1-139-02041-2 (ebook). xviii + 643 pp. LCCN QA188 .H66 2012.

Hoffman:1961:LA

[HK61] Kenneth Hoffman and Ray Alden Kunze. Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1961. 332 pp. LCCN QA251 .H67 1961.

Hoffman:1971:LA

[HK71] Kenneth Hoffman and Ray Alden Kunze. Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1971. ISBN 0-13-536797-2. viii + 407 pp. LCCN QA251 .H67 1971.

Hardy:1934:I

[HLP34] Godfrey H. Hardy, John Edensor Littlewood, and George Pólya. Inequalities. Cambridge University Press, Cambridge, UK, 1934. xii + 314 pp. LCCN QA303 .H25.

Hardy:1959:I

[HLP59] Godfrey H. Hardy, John Edensor Littlewood, and George Pólya. Inequalities. Cambridge University Press, Cambridge, UK, second edition, 1959. xii + 324 pp. LCCN QA303 .H25.

Hogben:2007:HLA

[Hog07] Leslie Hogben, editor. Handbook of Linear Algebra. Discrete Mathematics and its Applications (Boca Raton). Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-510-6 (hardcover), 1-4200-1057-3 (e-book). xxx + 1370 pp. LCCN QA184.2 .H36 2007.

URL http://www.crcnetbase.com/isbn/9781420010572;   http:
//www.crcnetbase.com/isbn/9781584885108;   http://www.
loc.gov/catdir/enhancements/fy0647/2006045491-d.html. Asso-

ciate editors: Richard Brualdi, Anne Greenbaum and Roy Mathias.

Householder:1964:TMN

[Hou64] Alston S. Householder. The Theory of Matrices in Numerical Analysis. Blaisdell Publishing Co. Ginn and Co., New York / Toronto / London, 1964. xi + 257 pp. LCCN QA263 .H67.

Householder:1972:LNA

[Hou72] A. S. Householder. Lectures on Numerical Algebra. Mathematical Association of America, Washington, DC, USA, 1972. iii + 250 + vii pp. LCCN QA155 .H67. Notes on lectures given at the June 19–July 28, 1972, MAA Summer Seminar, Williams College, Williamstown, MA, USA.

Householder:1975:TMN

[Hou75] Alston Scott Householder. The Theory of Matrices in Numerical Analysis. Dover, New York, NY, USA, 1975. ISBN 0-486-61781-5. xi + 257 pp. LCCN QA188 .H67 1975.

Householder:2006:TMN

[Hou06] Alston Scott Householder. The Theory of Matrices in Numerical Analysis. Dover, New York, NY, USA, dover edition, 2006. ISBN 0-486-44972-6 (paperback). xi + 257 pp. LCCN QA188 .H67

2006. URL http://www.loc.gov/catdir/enhancements/fy0634/
2005054775-d.html.

Hirsch:1974:DED

[HS74] Morris W. Hirsch and Stephen Smale. Differential Equations, Dynamical Systems, and Linear Algebra, volume 60 of Pure and Applied Mathematics. Academic Press, New York, NY, USA, 1974. ISBN 0-12-349550-4. xi + 358 pp. LCCN QA3 .P8 vol. 60 QA372. URL http://

www.loc.gov/catdir/description/els031/73018951.html;
http://www.loc.gov/catdir/toc/els031/73018951.html.

Huppert:1996:HWM

[HS96] Bertram Huppert and Hans Schneider, editors. Helmut Wielandt: Mathematische werke = Mathematical Works. Linear algebra and analysis, volume 2. Walter de Gruyter, New York, NY, USA, 1996. ISBN 3-11-012453-X (hardcover vol. 2), 3-11-012452-1 (vol. 2). xx + 632 pp. LCCN QA3 .W52 1996.

Jacobson:1943:TR

[Jac43] Nathan Jacobson. The Theory of Rings, volume II of American Mathematical Society Mathematical Surveys. American Mathematical Society, Providence, RI, USA, 1943. vi + 150 pp. LCCN QA247.

Kaplansky:2003:LAG

[Kap03] Irving Kaplansky. Linear Algebra and Geometry: a Second Course. Dover, New York, NY, USA, revised edition, 2003. ISBN 0-8284-0279-5 (hardcover), 0-486-43233-5 (paperback). xiv + 143 pp.

LCCN QA184.2 .K37 2003. URL http://www.loc.gov/catdir/
enhancements/fy0614/2003048902-d.html.

Karlin:1968:TP

[Kar68] Samuel Karlin. Total Positivity. Stanford University Press, Stanford, CA, USA, 1968. xi + 576 pp. LCCN QA331.5 .K3.

Kato:1966:PTL

[Kat66] Tosio Katō. Perturbation Theory for Linear Operators, volume 132 of Die Grundlehren der mathematischen Wissenschaften in Einseldarstellungen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. xix + 592 pp. LCCN QA320 .K33.

Kato:1976:PTL

[Kat76] Tosio Katō. Perturbation Theory for Linear Operators, volume 132 of Grundlehren der Mathematischen Wissenschaften. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1976. ISBN 0-387-07558-5. xxi + 619 pp. LCCN QA329.2 .K37 1976.

Kato:1984:PTL

[Kat84] Tosio Katō. Perturbation Theory for Linear Operators, volume 132 of Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second [corrected] edition, 1984. ISBN 3-540-07558-5 (Berlin), 0-387-07558-5 (New York). xxi + 619 pp. LCCN QA329.2 .K37 1984.

Kato:1995:PTL

[Kat95] Tosio Katō. Perturbation Theory for Linear Operators. Classics in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58661-X. xxi + 619

pp. LCCN QA329.2 .K37 1995. URL http://www.loc.gov/catdir/
enhancements/fy0815/94039131-b.html.

Kellogg:1971:TMT

[Kel71] R. Bruce Kellogg. Topics in matrix theory. Lecture notes 71.04, Chalmers Institute of Technology and the University of Göteborg, Göteborg, Sweden, 1971. 76 pp.

Kowalsky:1995:LAG

[KM95] Hans-Joachim Kowalsky and Gerhard O. Michler. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, tenth edition, 1995. ISBN 3-11-014501-4 (paperback), 3-11-014502-2 (hardcover). xiv + 399 pp. LCCN QA184 .K72 1995. URL http:

//www.gbv.de/dms/hbz/toc/ht006543541.pdf;   http://www.
zentralblattmath.org/zmath/en/search/?an=0827.15001.

Kowalsky:1998:LAG

[KM98] Hans-Joachim Kowalsky and Gerhard O. Michler. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, eleventh edition, 1998. ISBN 3-11-016185-0 (paperback), 3-11-016186-9 (hardcover). xiv + 401 pp. LCCN QA184 .K72 1998. URL http://

www.gbv.de/dms/hbz/toc/ht008883781.pdf.

Kowalsky:2003:LAG

[KM03] Hans-Joachim Kowalsky and Gerhard O. Michler. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, twelfth edition, 2003. ISBN 3-11-017963-6 (paperback), 3-11-020004-X (e-book). xv + 416 pp. LCCN QA184

.K693 2003. URL http://d-nb.info/968354416/04;   http://
www.degruyter.com/doi/book/10.1515/9783110200041;
http://www.degruyter.com/viewbooktoc/product/36737;
http://www.reference-global.com/isbn/978-3-11-020004-1.

Konig:1936:TEU

[Kőn36] Dénes Kőnig. Theorie der endlichen und unendlichen Graphen: kombinatorische Topologie der Streckenkomplexe. (German) [Theory of finite and infinite graphs: combinatorial topology of route [??] complexes], volume 16 of Mathematik und ihre Anwendungen in Monographien und Lehrbūchern. Teubner, Stuttgart, Germany; Leipzig, Germany, 1936. xi + 258 pp. LCCN QA611 .K6.

Konig:1950:TEU

[Kőn50] Dénes Kőnig. Theorie der endlichen und unendlichen Graphen. (German) [Theory of finite and infinite graphs]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1950. 348 pp. LCCN QA611 .K6 1950.

Konig:1954:TEU

[Kőn54] Dénes Kőnig. Theorie der endlichen und unendlichen Graphen. (German) [Theory of finite and infinite graphs]. Teubner, Stuttgart, Germany; Leipzig, Germany, 1954. 348 pp. LCCN QA611 .K6 1954.

Konig:1990:TFI

[Kőn90] Dénes Kőnig. Theory of Finite and Infinite Graphs. Birkhäuser Boston Inc., Cambridge, MA, USA, 1990. ISBN 3-7643-3389-8 (Basel), 0-8176-3389-8 (Boston), 1-4684-8971-2 (e-book). vi + 426 pp. LCCN QA166

.K6613 1990. URL http://www.gbv.de/dms/hbz/toc/ht003460259.
pdf;   http://www.zentralblatt-math.org/zmath/en/search/?

an=0695.05015. Translated from the German by Richard McCoart. With a commentary by W. T. Tutte and a biographical sketch by T. Gallai.

Kowalsky:1963:LAG

[Kow63] Hans-Joachim Kowalsky. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, 1963. 340 pp. LCCN ????

Kowalsky:1967:LAG

[Kow67] Hans-Joachim Kowalsky. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, third edition, 1967. 342 pp. LCCN QA251 .K68 1967.

Kowalsky:1969:LAG

[Kow69] Hans-Joachim Kowalsky. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, fourth edition, 1969. 342 pp. LCCN ????

Kowalsky:1970:LAG

[Kow70] Hans-Joachim Kowalsky. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, fifth edition, 1970. 341 pp. LCCN QA251 .K68 1970.

Kowalsky:1979:LAG

[Kow79] Hans-Joachim Kowalsky. Lineare Algebra. (German) [Linear Algebra]. Walter de Gruyter, New York, NY, USA, ninth edition, 1979. ISBN 3-11-008164-4, 3-11-007835-X. 367 pp. LCCN QH 140.

Kreyszig:1978:IFA

[Kre78] Erwin Kreyszig. Introductory Functional Analysis with Applications. Wiley, New York, NY, USA, 1978. ISBN 0-471-50731-8 (hardcover), 0-471-50459-9 (paperback). xiv + 688 pp. LCCN QA320

.K74. URL http://www.gbv.de/dms/bowker/toc/9780471504597.
pdf;   http://www.zentralblattmath.org/zmath/en/search/?
an=0706.46001.

Lancaster:1969:TM

[Lan69] Peter Lancaster. Theory of Matrices. Academic Press, New York, NY, USA, 1969. xii + 316 pp. LCCN QA263 .L33.

Lang:1987:LA

[Lan87] Serge Lang. Linear Algebra. Undergraduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1987. ISBN 0-387-96412-6. x + 285

pp. LCCN QA251 .L26 1987. URL http://www.loc.gov/catdir/
enhancements/fy0814/86021943-d.html;   http://www.loc.gov/
catdir/enhancements/fy0814/86021943-t.html.

Lax:2007:LAA

[Lax07] Peter D. Lax. Linear Algebra and its Applications. Pure and Applied Mathematics (Hoboken). Wiley-Interscience, New York, NY, USA, second edition, 2007. ISBN 0-471-75156-1 (hardcover). xvi + 376

pp. LCCN QA184.2 .L38 2007. URL http://www.loc.gov/catdir/
enhancements/fy0826/2007023226-b.html;   http://www.loc.
gov/catdir/enhancements/fy0826/2007023226-d.html;
http://www.loc.gov/catdir/toc/ecip0719/2007023226.html.

Lay:1982:CST

[Lay82] Steven R. Lay. Convex Sets and Their Applications. Pure and applied mathematics. Wiley, New York, NY, USA, 1982. ISBN 0-471-09584-2. ISSN 0079-8185. xvi + 244 pp. LCCN QA640 .L38 1982.

Lay:1992:CST

[Lay92] Steven R. Lay. Convex Sets and Their Applications. Krieger Pub. Co., Malabar, FL, USA, 1992. ISBN 0-89464-537-4. xvi + 244 pp. LCCN QA640 .L38 1991.

Lay:2007:CST

[Lay07] Steven R. Lay. Convex Sets and Their Applications. Dover, New York, NY, USA, dover edition, 2007. ISBN 0-486-45803-2 (paperback). xii + 244 pp. LCCN QA640 .L38 2007. URL http://www.loc.gov/catdir/

enhancements/fy0707/2007000125-d.html.

Lawson:1974:SLS

[LH74] Charles L. Lawson and Richard J. Hanson. Solving Least Squares Problems. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1974. ISBN 0-13-822585-0. xii + 340 pp. LCCN QA275 .L38.

Lawson:1995:SLS

[LH95] Charles L. Lawson and Richard J. Hanson. Solving Least Squares Problems, volume 15 of Classics in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995. ISBN 0-89871-356-0 (paperback). xii + 337 pp. LCCN QA275 .L38

1995. URL http://www.loc.gov/catdir/enhancements/fy0617/
95035178-d.html;   http://www.loc.gov/catdir/enhancements/
fy0617/95035178-t.html.

Lancaster:1985:TMA

[LT85] Peter Lancaster and Miron Tismenetsky. The Theory of Matrices: with Applications. Computer Science and Applied Mathematics. Academic Press, New York, NY, USA, second edition, 1985. ISBN 0-12-435560-9. xv + 570 pp. LCCN QA188 .L36 1985. URL http://www.loc.gov/

catdir/description/els032/83015775.html;
http://www.loc.gov/catdir/toc/els031/83015775.html.

MacDuffee:1933:TM

[Mac33] C. C. (Cyrus Colton) MacDuffee. The Theory of Matrices, volume 5 of Ergebnisse der Mathematik und ihrer Grenzgebiete … 2. J. Springer, Berlin, Germany, 1933. v + 110 pp. LCCN QA263 .M32.

MacDuffee:1946:TM

[Mac46] C. C. (Cyrus Colton) MacDuffee. The Theory of Matrices. Chelsea, New York, NY, USA, second edition, 1946. v + 110 pp. LCCN QA263 .M32.

MacDuffee:2004:TM

[Mac04] Cyrus Colton MacDuffee. The Theory of Matrices. Dover phoenix editions. Dover, New York, NY, USA, 2004. ISBN 0-486-49590-6. x + 110 pp. LCCN QA188 .M33 2004. URL http://www.loc.gov/

catdir/enhancements/fy0617/2004045572-d.html.

Marcus:1973:FDM

[Mar73] Marvin Marcus. Finite Dimensional Multilinear Algebra, volume 23(1) of Pure and applied mathematics. Marcel Dekker, Inc., New York, NY, USA, 1973. ISBN 0-8247-6077-8 (vol. 1). x + 292 pp. LCCN QA184 .M37.

Marcus:1975:FDM

[Mar75] Marvin Marcus. Finite Dimensional Multilinear Algebra, volume 23(2) of Pure and applied mathematics. Marcel Dekker, Inc., New York, NY, USA, 1975. ISBN 0-8247-6203-7 (vol. 2). xvi + 718 pp. LCCN QA184 .M37.

Minc:1988:NM

[Min88] Henryk Minc. Nonnegative Matrices. Wiley-Interscience Series in Discrete Mathematics and Optimization. Wiley, New York, NY, USA, 1988. ISBN 0-471-83966-3. xiii + 206 pp. LCCN QA188 .M558

1988. URL http://www.loc.gov/catdir/description/wiley034/
87027416.html;
http://www.loc.gov/catdir/toc/onix05/87027416.html.

Mirsky:1955:ILA

[Mir55] Leonid Mirsky. An Introduction to Linear Algebra. Clarendon Press, Oxford, UK and New York, NY, USA, 1955. xi + 433 pp. LCCN QA251 .M5.

Mirsky:1961:ILA

[Mir61] Leonid Mirsky. An Introduction to Linear Algebra. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1961. xi + 440 pp. LCCN QA251 .M5. Reprinted in [Mir90].

Mirsky:1963:ILA

[Mir63] Leonid Mirsky. An Introduction to Linear Algebra. Clarendon Press, Oxford, UK and New York, NY, USA, 1963. xi + 433 pp. LCCN ????

Mirsky:1982:ILA

[Mir82] Leonid Mirsky. An Introduction to Linear Algebra. Dover, New York, NY, USA, 1982. ISBN 0-486-61547-2. xi + 433 pp. LCCN QA184 .M57

1982. URL http://www.loc.gov/catdir/description/dover033/
90040299.html.

Mirsky:1990:ILA

[Mir90] Leonid Mirsky. An Introduction to Linear Algebra. Dover, New York, NY, USA, 1990. ISBN 0-486-66434-1 (paperback). viii + 440 pp. LCCN

QA184 .M57 1990. URL http://www.loc.gov/catdir/description/
dover033/90040299.html;   http://www.zentralblatt-math.
org/zmath/en/search/?an=0766.15001.

Marcus:1964:SMT

[MM64] Marvin Marcus and Henryk Minc. A Survey of Matrix Theory and Matrix Inequalities. Allyn and Bacon, Needham Heights, MA, USA, 1964. xvi + 180 pp. LCCN QA263 .M345.

Marcus:1992:SMT

[MM92] Marvin Marcus and Henryk Minc. A Survey of Matrix Theory and Matrix Inequalities. Dover, New York, NY, USA, 1992. ISBN 0-486-67102-X (paperback). xii + 180 pp. LCCN QA188 .M36 1992. URL http:/

/www.loc.gov/catdir/description/dover032/91039704.html.

Marshall:1979:ITM

[MO79] Albert W. Marshall and Ingram Olkin. Inequalities: Theory of Majorization and Its Applications, volume 143 of Mathematics in Science and Engineering. Academic Press, New York, NY, USA, 1979. ISBN 0-12-473750-1. xx + 569 pp. LCCN QA295 .M42. URL http://

www.loc.gov/catdir/description/els031/79050218.html;
http://www.loc.gov/catdir/toc/els031/79050218.html.

Marshall:2011:ITM

[MOA11] Albert W. Marshall, Ingram Olkin, and Barry C. Arnold. Inequalities: Theory of Majorization and Its Applications, volume ?? of Springer Series in Statistics. Springer Science+Business Media, LLC, New York, NY, USA, second edition, 2011. CODEN ???? ISBN 0-387-40087-7, 0-387-68276-7 (e-book). ISSN 0172-7397. xxvii + 909 pp. LCCN QA295 .M37 2011; QA295 .M42

2011. URL http://link.springer.com/book/10.1007/978-0-387-
68276-1;   http://www.loc.gov/catdir/enhancements/fy1114/
2010931704-b.html;   http://www.loc.gov/catdir/
enhancements/fy1114/2010931704-d.html;   http://www.loc.
gov/catdir/enhancements/fy1114/2010931704-t.html.

Muir:1906:TDH

[Mui23] Sir Thomas Muir. The Theory of Determinants in the Historical Order of Development. Macmillan Publishing Company, New York, NY, USA and London, UK, 1906, 1911, 1920, 1923. xii + 491 (vol. 1), xvi + 475 (vol. 2), xxvi + 503 (vol. 3), xxxi + 508 (vol. 4) pp. LCCN QA191 .M9. Four volumes.

Muir:1930:CHD

[Mui30] Sir Thomas Muir. Contributions to the History of Determinants, 1900– 1920. Blackie and Son, London and Glasgow, UK, 1930. xxiii + 1 + 408 pp. LCCN QA191 .M88.

Muir:1960:TDH

[Mui60a] Sir Thomas Muir. The Theory of Determinants in the Historical Order of Development. Dover, New York, NY, USA, 1960. ???? pp. LCCN QA191 .M92. 4 original volumes in 2 reprint volumes.

Muir:1960:TTD

[Mui60b] Sir Thomas Muir. A Treatise on the Theory of Determinants. Dover, New York, NY, USA, 1960. 766 pp. LCCN QA191 .M85 1960. Corrected printing of the 1933 edition, revised and enlarged by William H. Metzler.

Noble:1969:ALA

[ND69] Ben Noble and James W. Daniel. Applied Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1969. xvi + 523 pp. LCCN QA184 .N6 1969.

Noble:1977:ALA

[ND77] Ben Noble and James W. Daniel. Applied Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1977. ISBN 0-13-041343-7. xvii + 477 pp. LCCN QA184 .N6 1977.

Noble:1988:ALA

[ND88] Ben Noble and James W. Daniel. Applied Linear Algebra. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1988. ISBN 0-13-041260-0, 0-13-040957-X (paperback). xvi + 521 pp.

LCCN QA184 .N61 1988. URL http://www.gbv.de/dms/bowker/
toc/9780130412607.pdf.

Noble:1989:ALA

[NDG89] Ben Noble, James W. Daniel, and Virgilio González Pozo. Algebra lineal aplicada. Prentice Hall Hispanoamericana, México, DF, México, 1989. ISBN 968-880-173-9, 0-13-041260-0. xvii + 572 pp. LCCN QA184 .N618 1989.

Newman:1972:IM

[New72] Morris Newman. Integral Matrices, volume 45 of Pure and Applied Mathematics. Academic Press, New York, NY, USA, 1972. ISBN 0-12-517850-6. xvii + 224 pp. LCCN QA3 .P8 1972.

OMeara:2011:ATL

[OCV11] Kevin C. O’Meara, John Clark, and Charles I. Vinsonhaler. Advanced Topics in Linear Algebra: Weaving Matrix Problems Through the Weyr Form. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-979373-5 (hardcover). xxii + 400 pp. LCCN QA184.2 .O44 2011.

Ortega:1987:MTS

[Ort87] James McDonough Ortega. Matrix Theory: a Second Course. The University Series in Mathematics. Plenum Press, New York, NY, USA; London, UK, 1987. ISBN 0-306-42433-9, 1-4899-0471-9 (e-book). xii + 262 pp. LCCN QA188 .O78 1987. URL http://www.gbv.de/dms/

hbz/toc/ht002910694.pdf;   http://www.zentralblatt-math.
org/zmath/en/search/?an=0654.15001.

Parlett:1980:SEP

[Par80] Beresford N. Parlett. The Symmetric Eigenvalue Problem. Series in Computational Mathematics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-880047-2. xix + 348 pp. LCCN QA188 .P3. Prentice-Hall Series in Computational Mathematics.

Parlett:1998:SEP

[Par98] Beresford N. Parlett. The Symmetric Eigenvalue Problem, volume 20 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-402-8 (paperback), 1-61197-116-0 (e-book). xxiv + 398 pp. LCCN QA188 .P37

1998. URL http://www.loc.gov/catdir/enhancements/fy0708/

97040623-d.html. Corrected reprint [Par80].

Parshall:2006:JJS

[Par06] Karen Hunger Parshall. James Joseph Sylvester: Jewish Mathematician in a Victorian World. The Johns Hopkins University Press, Baltimore, MD, USA, 2006. ISBN 0-8018-8291-5. xiii + 461 pp.

LCCN QA29.S96 P375 2006. URL http://www.loc.gov/catdir/
enhancements/fy0625/2005052035-b.html;   http://www.loc.
gov/catdir/enhancements/fy0625/2005052035-d.html;
http://www.loc.gov/catdir/toc/fy0701/2005052035.html.

Perlis:1952:TM

[Per52] Sam Perlis. Theory of Matrices. Addison-Wesley, Reading, MA, USA, 1952. xiv + 237 pp. LCCN QA263 .P4.

Perlis:1991:TM

[Per91] Sam Perlis. Theory of Matrices. Dover, New York, NY, USA, 1991. ISBN 0-486-66810-X (paperback). xiv + 237 pp. LCCN QA188 .P42

1991. URL http://www.loc.gov/catdir/description/dover031/
91019037.html.

Pringle:1971:GIMa

[PR71a] R. M. Pringle and Arthur Asquith Rayner. Generalized Inverse Matrices with Applications to Statistics, volume 28 of Griffin’s statistical monographs and courses. Griffin, London, UK, 1971. ISBN 0-85264-181-8. viii + 127 pp. LCCN QA279 .P74 1971b.

Pringle:1971:GIMb

[PR71b] R. M. Pringle and Arthur Asquith Rayner. Generalized Inverse Matrices with Applications to Statistics, volume 28 of Griffin’s statistical monographs and courses. Hafner Pub. Co., New York, NY, USA, 1971. viii + 127 pp. LCCN QA279 .P74. URL http://www.gbv.de/dms/

hbz/toc/ht000763299.pdf.

Rogers:1980:MD

[Rog80] Gerald Stanley Rogers. Matrix Derivatives, volume 2 of Lecture Notes in Statistics. Marcel Dekker, Inc., New York, NY, USA, 1980. ISBN 0-8247-1176-9 (paperback). v + 209 pp. LCCN QA188 .R63. URL

http://www.gbv.de/dms/hbz/toc/ht002508073.pdf.

Roman:1992:ALA

[Rom92] Steven Roman. Advanced Linear Algebra, volume 135 of Graduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 1-4757-2180-3, 1-4757-2178-1 (e-book). ISSN 0072-5285. xii + 370 pp. LCCN QA184-205.

Roman:2005:ALA

[Rom05] Steven Roman. Advanced Linear Algebra, volume 135 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005. ISBN 0-387-27474-X, 0-387-24766-1. ISSN 0072-5285. xvi + 482 pp. LCCN QA184.2 .R66 2005.

Roman:2008:ALA

[Rom08] Steven Roman. Advanced Linear Algebra, volume 135 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 0-387-72828-7, 0-387-72831-7. ISSN 0072-5285. xviii + 522 pp. LCCN QA184 .R66

2008; QA184.2 .R66 2007. URL http://www.gbv.de/dms/bowker/
toc/9780387728285.pdf;   http://www.zentralblatt-math.org/
zmath/en/search/?an=1132.15002.

Radjavi:2000:ST

[RR00] Heydar Radjavi and Peter Rosenthal. Simultaneous Triangularization. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98467-4 (hardcover), 0-387-98466-6 (paperback), 1-4612-1200-6 (e-book). xii + 318

pp. LCCN QA197 .R33 2000. URL http://www.loc.gov/catdir/
enhancements/fy0815/99023772-d.html;   http://www.loc.gov/
catdir/enhancements/fy0815/99023772-t.html.

Rudin:1953:PMA

[Rud53] Walter Rudin. Principles of Mathematical Analysis. International series in pure and applied mathematics. McGraw-Hill, New York, NY, USA, 1953. 227 pp. LCCN QA300 .R8.

Rudin:1964:PMA

[Rud64] Walter Rudin. Principles of Mathematical Analysis. International series in pure and applied mathematics. McGraw-Hill, New York, NY, USA, second edition, 1964. ix + 270 pp. LCCN QA300 .R8 1964.

Rudin:1976:PMA

[Rud76] Walter Rudin. Principles of Mathematical Analysis. International Series in Pure and Applied Mathematics. McGraw-Hill, New York, NY, USA, third edition, 1976. ISBN 0-07-054235-X. x + 342 pp. LCCN

QA300 .R8 1976. URL http://www.loc.gov/catdir/enhancements/
fy0602/75017903-d.html;
http://www.loc.gov/catdir/toc/mh031/75017903.html.

Seneta:1973:NNMc

[Sen73a] E. (Eugene) Seneta. Non-Negative Matrices: an Introduction to Theory and Applications. Allen and Unwin, London, UK, 1973. ISBN 0-04-519011-9. x + 214 pp. LCCN QA188 .S46 1973b.

Seneta:1973:NNMa

[Sen73b] Eugene Seneta. Non-Negative Matrices: an Introduction to Theory and Applications. Halsted Press, New York, USA, 1973. ISBN 0-470-77605-6. x + 214 pp. LCCN QA188 .S46.

Seneta:1981:NNM

[Sen81] Eugene (Eugene) Seneta. Non-Negative Matrices and Markov Chains. Springer Series in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-90598-7 (New York), 3-540-90598-7 (Berlin). xv + 279 pp. LCCN QA188 .S46

1981. URL http://www.gbv.de/dms/bowker/toc/9783540905981.
pdf.

Seneta:2006:NNM

[Sen06] E. (Eugene) Seneta. Non-Negative Matrices and Markov Chains. Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., revised edition, 2006. ISBN 0-387-29765-0 (paperback). xv + 279 + 8 pp. LCCN QA188 .S46

2006. URL http://www.loc.gov/catdir/enhancements/fy0663/
2005935207-d.html;
http://www.loc.gov/catdir/toc/fy0804/2005935207.html.

Serre:2002:MTA

[Ser02] Dénis Serre. Matrices: Theory and Applications, volume 216 of Graduate Texts in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95460-0, 0-387-22758-X (e-book). xvi + 202 pp. LCCN QA188 .S4713

2002. URL http://link.springer.com/book/10.1007/b98899;
http://www.loc.gov/catdir/enhancements/fy0817/2002022926-
b.html;   http://www.loc.gov/catdir/enhancements/fy0817/
2002022926-d.html;   http://www.loc.gov/catdir/
enhancements/fy0817/2002022926-t.html;
http://www.springerlink.com/openurl.asp?genre=book&isbn=

978-0-387-95460-8. Translated from the 2001 French original.

Stewart:1990:MPT

[SS90] Gilbert Wright (Gilbert W.) Stewart and Ji-guang Sun. Matrix Perturbation Theory. Computer science and scientific computing. Academic Press, New York, NY, USA, 1990. ISBN 0-12-670230-6 (hardcover). xv + 365 pp. LCCN QA871 .S775 1990. URL http://www.loc.gov/

catdir/description/els032/90033378.htm;
http://www.loc.gov/catdir/description/els032/90033378.
html;   http://www.loc.gov/catdir/toc/els031/90033378.htm;
 http://www.loc.gov/catdir/toc/els031/90033378.html.

Suprunenko:1968:CM

[ST68] D. A. (Dmitrij Alekseevic) Suprunenko and R. I. (Regina Iosifovna) Tyshkevich. Commutative Matrices. Academic Press, New York, NY, USA, 1968. viii + 158 pp. LCCN QA263 .S913.

Staib:1969:IML

[Sta69] John H. Staib. An Introduction to Matrices and Linear Transformations. Addison-Wesley series in mathematics. Addison-Wesley, Reading, MA, USA, 1969. xii + 336 pp. LCCN QA251 .S77.

Stewart:1973:IMC

[Ste73] G. W. (Gilbert W.) Stewart. Introduction to Matrix Computations. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, NY, USA, 1973. ISBN 0-12-670350-7. xiii + 441 pp. LCCN QA188 .S71 1973. URL http://catdir.

loc.gov/catdir/description/els031/72082636.html;

http://catdir.loc.gov/catdir/toc/els031/72082636.html;
http://www.gbv.de/dms/hbz/toc/ht000832348.pdf;   http://
www.loc.gov/catdir/description/els031/72082636.html;
http://www.loc.gov/catdir/toc/els031/72082636.html.

Strutt:1877:TS

[Str78] John William Strutt (Baron Rayleigh). The Theory of Sound. Macmillan Publishing Company, New York, NY, USA and London, UK, 1877– 1878. xi + 326 (vol. 1), x + 302 (vol. 2) pp. LCCN QC223 .R26.

Strutt:1894:TS

[Str96] John William Strutt (Baron Rayleigh). The Theory of Sound. Macmillan Publishing Company, New York, NY, USA and London, UK, second edition, 1894–1896. xiv + 480 (vol. 1), xvi + 504 (vol. 2) pp. LCCN QC223 .R26.

Strutt:1945:TS

[Str45] John William Strutt (Baron Rayleigh). The Theory of Sound. Dover, New York, NY, USA, second edition, 1945. xlii + 480 (vol. 1), xii + 504 (vol. 2) pp. LCCN QC223 .R26.

Turnbull:1932:ITC

[TA32] Herbert Westren Turnbull and Alexander Craig Aitken. An Introduction to the Theory of Canonical Matrices. Blackie, Glasgow, Scotland, 1932. 192 pp. LCCN QA263 .T8.

Turnbull:1952:ITC

[TA52] Herbert Westren Turnbull and Alexander Craig Aitken. An Introduction to the Theory of Canonical Matrices. Blackie, Glasgow, Scotland, second edition, 1952. xiii + 200 pp. LCCN ????

Turnbull:1961:ITC

[TA61] H. W. Turnbull and A. C. Aitken. An Introduction to the Theory of Canonical Matrices. Dover, New York, NY, USA, third edition, 1961. xiii + 200 pp. LCCN QA263 .T8 1961. Reprint of [TA52].

Turnbull:2005:ITC

[TA05] H. W. (Herbert Westren) Turnbull and A. C. (Alexander Craig) Aitken. Introduction to the Theory of Canonical Matrices. Dover, New York, NY, USA, 2005. ISBN 0-486-44168-7. xiii + 200 pp. LCCN

QA263 .T8 2005. URL http://www.loc.gov/catdir/enhancements/
fy0619/2004061766-d.html;   http://www.loc.gov/catdir/
enhancements/fy1318/2004061766-t.html.

Tang:2007:MMEa

[Tan07a] Kwong-Tin Tang. Mathematical Methods for Engineers and Scientists: 1. Complex Analysis, Determinants and Matrices. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-30273-5 (hardcover), 3-540-30274-3 (e-book). x + 319 pp.

LCCN QA401 .T36 2007. URL http://d-nb.info/97689131X/04;
 http://swbplus.bsz-bw.de/bsz261984101cov.htm;
http://swbplus.bsz-bw.de/bsz261984101inh.pdf;
http://swbplus.bsz-bw.de/bsz261984101kap.pdf;   http://
swbplus.bsz-bw.de/bsz261984101vlg.htm;   http://www.loc.
gov/catdir/enhancements/fy0824/2006932619-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy0824/2006932619-d.html;
 http://www.loc.gov/catdir/toc/fy0804/2006932619.html.

Tang:2007:MMEb

[Tan07b] Kwong-Tin Tang. Mathematical Methods for Engineers and Scientists: 2. Vector Analysis, Ordinary Differential Equations and Laplace Transforms. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-30268-9 (hardcover), 3-540-30270-0 (e-book). xi + 339 pp. LCCN QA401 .T36 2007. URL http://d-nb.

info/976891336/04.

Tang:2007:MMEc

[Tan07c] Kwong-Tin Tang. Mathematical Methods for Engineers and Scientists: 3. Fourier Analysis, Partial Differential Equations and Variational Methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-44695-8 (hardcover), 3-540-44697-4 (e-book). xi + 438 pp. LCCN QA401 .T36 2007.

Todd:1962:SNA

[Tod62] John Todd, editor. Survey of Numerical Analysis. McGraw-Hill, New York, NY, USA, 1962. xvi + 589 pp. LCCN QA297 .T6. URL http:

//en.wikipedia.org/wiki/John_Todd_(computer_scientist).

Turnbull:1950:TDM

[Tur50] H. W. (Herbert Westren) Turnbull. Theory of Determinants, Matrices and Invariants. Blackie, Glasgow, Scotland, second edition, 1950. xvi + 368 pp. LCCN ????

Valentine:1964:CS

[Val64] Frederick A. Valentine. Convex Sets. McGraw-Hill Series in Higher Mathematics. McGraw-Hill, New York, NY, USA, 1964. ix + 238 pp. LCCN QA248 .V3.

Valentine:1976:CS

[Val76] Frederick A. (Frederick Albert) Valentine. Convex Sets. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1976. ISBN 0-88275-289-8. ix + 238 pp. LCCN QA640 .V34 1975.

Varga:2000:MIA

[Var00] Richard S. Varga. Matrix Iterative Analysis, volume 27 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 3-642-05154-5 (paperback), 3-540-66321-5 (hardcover), 3-642-05156-1 (e-book). x + 358 pp. LCCN QA297.8 .V37

2000. URL http://lcweb.loc.gov/catdir/toc/99050196.html;
http://www.loc.gov/catdir/enhancements/fy0815/99050196-d.
html;   http://www.springerlink.com/openurl.asp?genre=
book&isbn=978-3-642-05161-6.

Varga:2004:GHC

[Var04] Richard S. Varga. Geršgorin and His Circles, volume 36 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21100-4 (hardcover), 3-642-17798-0 (e-book). ISSN 0179-3632. x + 226 pp. LCCN QA193 .V37 2004. URL http://www.loc.gov/

catdir/enhancements/fy0815/2004104814-d.html.

Wedderburn:1934:LM

[Wed34] J. H. M. (Joseph Henry Maclagan) Wedderburn. Lectures on Matrices, volume XVII of American Mathematical Society. Colloquium publications. American Mathematical Society, Providence, RI, USA, 1934. vii + 200 pp. LCCN QA1 .A5225 vol. 17a.

Wedderburn:1964:LM

[Wed64] J. H. M. (Joseph Henry Maclagan) Wedderburn. Lectures on Matrices, volume XVII of American Mathematical Society. Colloquium publications. Dover, New York, NY, USA, 1964. vii + 200 pp. LCCN QA1 .A5225 vol. 17a.

Wedderburn:2005:LM

[Wed05] J. H. M. (Joseph Henry Maclagan) Wedderburn. Lectures on Matrices. Dover phoenix editions. Dover, New York, NY, USA, 2005. ISBN 0-486-44167-9. vii + 200 pp. LCCN QA188 .W43 2005. URL http://

www.loc.gov/catdir/enhancements/fy0618/2004059120-d.html.

Wilkinson:1965:AEP

[Wil65] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Clarendon Press, Oxford, UK and New York, NY, USA, 1965. xviii + 662 pp. LCCN QA218 .W5.

Wilkinson:1988:AEP

[Wil88] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Monographs on Numerical Analysis. Clarendon Press, Oxford, UK and New York, NY, USA, 1988. ISBN 0-19-853418-3. xviii + 662 pp. LCCN QA218 .W5

1988. URL http://www.gbv.de/dms/bowker/toc/9780198534181.
pdf;   http://www.loc.gov/catdir/enhancements/fy0604/
87028274-d.html;   http://www.loc.gov/catdir/enhancements/

fy0604/87028274-t.html. Oxford Science Publications.

Wielandt:1967:TAT

[WM67] Helmut W. Wielandt and Robert R. Meyer. Topics in the Analytic Theory of Matrices: Lecture Notes. Department of Mathematics, University of Wisconsin, Madison, WI, USA, 1967. 129 pp. LCCN QA262 .F72; QA263 .W43. Reprinted in [HS96, 271–352].

Young:1972:SNM

[YG72] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics. Addison-Wesley, Reading, MA, USA, 1972. ISBN 0-201-08773-1, 0-486-65691-8 (Dover paperback). x + 492 (A1–A18 and B1–B14 and I1–I19) pp. LCCN QA297 .Y63 1972.

Young:1973:SNM

[YG73] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics. Addison-Wesley, Reading, MA, USA, 1973. 493–1099 + xviii + xvi + xviii pp. LCCN QA297 .Y63.

Young:1988:SNMa

[YG88a] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics, volume I. Dover, New York, NY, USA, 1988. ISBN 0-486-65691-8. x + 492 + A22 + B16 + I18 pp. LCCN QA297.Y63

1972. URL http://www.zentralblatt-math.org/zmath/en/search/

?an=0732.65002. Corrected reprint of the 1973 original.

Young:1988:SNMb

[YG88b] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics, volume II. Dover, New York, NY, USA, 1988. ISBN 0-486-65692-6. xii + 607 + A18 + B16 + I18 pp. LCCN QA297.Y63

1972. URL http://www.zentralblatt-math.org/zmath/en/search/

?an=0732.65003. Corrected reprint of the 1973 original.

Zurmuhl:1984:MIA

[ZF84] Rudolf Zurmühl and Sigurd Falk. Matrizen und ihre Anwendungen für angewandte Mathematiker, Physiker und Ingenieure. Teil 1. Grundlagen. (German) [Matrices and their Uses for Applied Mathematicians, Physicists, and Engineers. Part 1. Foundations]. Grundlagen. [Foundations]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 1984. ISBN 0-387-03238-X, 0-387-12848-4 (New York), 3-540-12848-4 (Berlin). xiv + 342 pp. LCCN QA188 .Z87 1984.

Zurmuhl:1986:MIA

[ZF86] Rudolf Zurmühl and Sigurd Falk. Matrizen und ihre Anwendungen für angewandte Mathematiker, Physiker und Ingenieure. Teil 2. Numerische Methoden. (German) [Matrices and their Uses for Applied Mathematicians, Physicists, and Engineers. Part 2. Numerical Methods]. Numerische Methoden. [Numerical methods]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 1986. ISBN 0-387-15474-4 (New York hardcover), 3-540-15474-4 (Berlin hardcover). xvi + 476 pp. LCCN QA188 .Z87 1984.

Zhan:2002:MI

[Zha02] Xingzhi Zhan. Matrix Inequalities, volume 1790 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNMAA2. ISBN 3-540-43798-3 (print), 3-540-45421-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 116 pp. LCCN QA3 .L28 no. 1790.

URL http://link.springer.com/book/10.1007/b83956;   http:
//www.loc.gov/catdir/enhancements/fy0817/2002070800-d.
html;   http://www.loc.gov/catdir/enhancements/fy0817/
2002070800-t.html;   http://www.springerlink.com/content/
978-3-540-45421-2;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0075-8434&volume=1790.

Zhang:2005:SCA

[Zha05] Fuzhen Zhang, editor. The Schur Complement and its Applications, volume 4 of Numerical Methods and Algorithms. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-24271-6 (hardcover), 0-387-24273-2 (e-book). xvi + 295

pp. LCCN QA184.2 .S38 2005. URL http://www.loc.gov/catdir/

enhancements/fy0663/2005042750-d.html;   http://www.loc.
gov/catdir/enhancements/fy0821/2005042750-t.html.

Zhang:2009:LAC

[Zha09] Fuzhen Zhang. Linear Algebra: Challenging Problems for Students. Johns Hopkins Studies in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, second edition, 2009. ISBN 0-8018-9125-6 (hardcover), 0-8018-9126-4 (paperback). xviii + 245 pp. LCCN QA184.5 .Z48 2009.

Zhang:2011:MTB

[Zha11] Fuzhen Zhang. Matrix Theory: Basic Results and Techniques. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2011. ISBN 1-4614-1098-3 (paperback), 1-4614-1099-1 (e-book). xviii + 399 pp. LCCN QA188 .Z47

2011. URL http://www.loc.gov/catdir/enhancements/fy1406/
2011935372-b.html;   http://www.loc.gov/catdir/
enhancements/fy1406/2011935372-d.html;   http://www.loc.
gov/catdir/enhancements/fy1406/2011935372-t.html.

Zurmuhl:1950:MDI

[Zur50] Rudolf Zurmühl. Matrizen: Eine Darstellung für Ingenieure. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1950. ISBN 3-642-53289-6. xiv + 427 pp. LCCN

QA263 .Z8. URL http://d-nb.info/1033262757/34;   http://
www.springerlink.com/content/978-3-642-53289-4.

Zurmuhl:1958:MDI

[Zur58] Rudolf Zurmühl. Matrizen: Eine Darstellung für Ingenieure. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1958. ISBN 3-642-53291-

8.  xv + 467 pp.  LCCN QA188.Z8 1958; QA263 .Z8

1958.  URL http://d-nb.info/1028197780/3;   http://nbn-
resolving.de/urn:nbn:de:1111-201211261057;
http://www.springerlink.com/content/978-3-642-53291-7.

Zurmuhl:1961:MIT

[Zur61] Rudolf Zurmühl. Matrizen und ihre technischen Anwendungen. (German) [Matrices and their technical applications]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1961. 459 pp. LCCN QA263 .Z8 1961.

Zurmuhl:1964:MIT

[Zur64] Rudolf Zurmühl. Matrizen und ihre technischen Anwendungen. (German) [Matrices and their technical applications]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 1964. xii + 452 pp. LCCN QA263 .Z8 1964.